OpenStudy (anonymous):

rationalize the denominator 1/√5

OpenStudy (anonymous):

\[(1/\sqrt{5}) * (\sqrt{5}/\sqrt{5})\]

OpenStudy (anonymous):

i dont understand how to do it..

OpenStudy (anonymous):

notice we need to multiply by \[\sqrt{5}/\sqrt{5}\] so that we will no longer have a radical in the denominator. since \[\sqrt{5}/\sqrt{5}\] is equal to 1, we are not changing the value. We are just changing what it looks like :)

OpenStudy (anonymous):

To rationalize the denominator of a fraction: \[\frac{k}{\sqrt{a}}\] You multiply it by a form of '1' that will rid you of the square root in the denominator: \[\frac{\sqrt{a}}{\sqrt{a}} = 1\] \[\implies \frac{k}{\sqrt{a}} = \frac{k}{\sqrt{a}} \cdot \frac{\sqrt{a}}{\sqrt{a}} = \frac{k\sqrt{a}}{a}\]

OpenStudy (anonymous):

so it would be 1√5/5?

OpenStudy (anonymous):

Yes, but \[1\sqrt{5} = 1 \cdot \sqrt{5} = \sqrt{5}\]

OpenStudy (anonymous):

So no need to write it with the extra 1

OpenStudy (anonymous):

so this problem \[23\div \sqrt{23}\] would be \[23\sqrt{23}\div23\] but wouldnt that equal \[\sqrt{23}\]?

OpenStudy (anonymous):

Yes